Assume that a sample is used to estimate a population mean μμ. Find the 80% confidence interval for a sample of size 457 with a mean of 66.5 and a standard deviation of 13.4. Enter your answer as a tri-linear inequality accurate to 3 decimal places.
____ < μ < ____
Answer #1

Solution Given that ř = 66.5 8 = 13.4 At 80% confidence level o = 1- 80% = 1- 0.8 = 0.2 a 0.2 0.1 2 2 2012 - 20. = 1.282 Marg